OpenStudy (anonymous):

in a assortment of calculators 12 are graphing calculators, 16 are scientific, 7 are business. Find the probability of randomly selecting a business calculator, replacing it, then randomly selecting a graphing calculator.

jimthompson5910 (jim_thompson5910):

There are 12+16+7 = 35 calculators total P(business calc then graphing calc) = P(business calc)*P(graphing calc) P(business calc then graphing calc) = (7/35)*(12/35) P(business calc then graphing calc) = (7*12)/(35*35) P(business calc then graphing calc) = (84)/(1225) P(business calc then graphing calc) = 12/175

OpenStudy (anonymous):

?